# Deep Learning for Cross-Domain Data Fusion in Urban Computing: Taxonomy, Advances, and Outlook

### v1

@@ 5.6 Environment
Rapid urbanization may result in a potential threat to cities’ environment. For example, it is reported that ninety percent of air pollution is caused by urban transportation. Environment protection is an essential topic for urban computing. We have witnessed much research on the environment from different aspects of urban computing, such as air quality prediction Kök et al. (2017); Iskandaryan et al. (2020); Kang et al. (2018); Wang et al. (2017a); Liang et al. (2023), noise controlling Abbaspour et al. (2015); JEZDOVIÄ et al. (2018); Jezdović et al. (2021); Dutta et al. (2017); Liu et al. (2020b) and weather forecasting Zheng et al. (2015); Yonekura et al. (2018); Chen et al. (2022a); Song et al. (2019b); Ghoneim et al. (2017). Various data sources cloud provide information from various aspects for us to sense the environment. Further, the fusion those information from multiple sources was proven to be effective in understanding and predicting the variation of the environment in a city.
Zheng et al. (2015) conducted research on cities’ air quality based on multi-source big data. They proposed a multi-view hybrid model to predict future 48-hour air quality of a single station point by converging historical air quality data of one station, current meteorological data in the area, weather forecasting data as well as air quality data from other stations. The various meteorological factors and spatial relations between different stations are both considered by effectively fusing these data which significantly improved the capability and accuracy of fine-grain air quality prediction. Further, Ma et al. (2023a) utilized a graph neural network to model spatio-temporal correlations between different meteorological variables and different stations as graph neural networks have demonstrated excellent performances for modeling spatio-temporal information. This kind of research makes city-level fine-grained weather forecasting possible with multi-source spatio-temporal data.
@@ 5.7.1 Energy Efficient Urban Planning
Electric vehicles are regarded as a promising solution for green transportation and sustainable cities. With the gradual expansion of the market share of electrical vehicles, the demand for car charging facilities is increasing. One important issue in urban planning is the selection of charging station locations. Proper selection of these locations not only optimizes city management but also meets the daily commuting needs of users, reducing unnecessary energy consumption during the search for charging stations. The selection of charging station locations should take into account various factors, such as the economic situation of the region, the number of electric vehicles in circulation, residents’ travel patterns, road conditions, and traffic conditions. This necessitates the integration and analysis of multiple data sources to estimate the regional charging demand and identify suitable locations for charging stations. Tu et al. (2016) proposed a spatio-temporal demand coverage location model based on the taxi GPS data in Shenzhen, China as well as data from the charging stations. This model provided a charging station sitting approach by considering the waiting time which could hugely reduce energy consumption for charging. He et al. (2018) developed a location method with the consideration of cars’ driving range extracted from the GPS data. Battaïa et al. (2023) proposed a framework for charging situation sitting considering the real-life problem of charging infrastructure and sustainability.
@@ 5.7.2 Energy Efficient Urban Transportation
Transportation energy efficiency aims to recommend and navigate energy-efficient travel routes for every transportation participant. Nowadays, within the worldwide energy shortage, fuel prices have been continuously rising Zhao and Yu (2017). However, traffic congestion in 439 cities in the United States resulted in a total loss of 3.3 billion gallons of wasted fuel Schrank et al. (2019) in only one year. Traditional route planning studies primarily emphasized faster and more reliable routes, often overlooking the differences in energy consumption among different travel routes and modes for the same travel demand. For instance, in certain situations, taking a route through a congested city center may result in faster arrival times compared to taking a detour on a highway. Despite the first route being shorter in distance, the driver may experience delays due to frequent stops and slow speeds caused by city congestion.
In application, recommending energy-efficient routes for users is a highly complex task. Unlike traditional shortest path problems, it requires considering various factors such as vehicle efficiency, driving habits, road conditions, POIs, and more. The key to addressing this task lies in effectively integrating information from multiple sources. Wang et al. (2022) researched personalized fuel-efficient route recommendations based on history trajectories. The temporal information from the trajectories as well as the driving factors are input into a transformer model to compose fuel efficiency prediction on the target route. By combining with a genetic algorithm for the best recommendation, this research could effectively find the fuel-efficient routes on the real-world dataset. DeepFEC Elmi and Tan (2021) proposed a framework to predict road-level energy consumption through a fusion of contextual vehicle data such as type, weight and engine configuration, and spatio-temporal traffic data on roads. In this framework, a deep convolution residual model is designed to capture the spatial dynamic information from the data and a Bi-LSTM model is responsible for the temporal information. Oh et al. (2019) is the first large dataset that combines trajectories of personal cars with users’ information, making it possible to mine users’ driving behaviors. Based on this multi-source dataset, Lai et al. (2023) designed a meta-learning framework for predicting vehicle energy consumption in Ann Arbor, Michigan, USA that can be fine-tuned based on a user’s driving preference.
@@ 6 Challenges and Future Directions
While urban multi-modal research has made significant advancements in recent years, several challenging issues persist, highlighting directions for potential future research. We summarize these challenges and suggest potentially feasible research directions as follows:
@@
  • LLM-enhanced Application: Since the advent of GPT-4 OpenAI (2023) and Sora Brooks et al. (2024), the academic community has embarked on extensive research on the roles of LLM as predictors Lin and Chen (2023); Hu et al. (2023b), enhancers Yin et al. (2023a); Peng et al. (2023), controllers Foosherian et al. (2023); Shen et al. (2023b); Xu et al. (2023b), and evaluators Svikhnushina and Pu (2023); Lin and Chen (2023). However, the field of cross-domain data fusion in urban computing is still in its initial stage in terms of exploring how to apply LLMs effectively. For example, there exists research Roberts et al. (2023); Bubeck et al. (2023); Manvi et al. (2023) focusing on LLM as urban predictors but also as naive reasoners yet; whereas others Hu et al. (2023a); Kuckreja et al. (2023) attempt to leverage LLM as model backbone but such works are limited within remote sensing domain instead of general urban computing. Therefore, we eagerly anticipate that the research community will focus on the application of LLMs in urban computing, exploring their impact similar to their roles in NLP and other domains.
@@
  • Agent-based Simulation: Agent-based simulation aims to model the behavior of individual components or agents to comprehend their interactions and how they collectively contribute to the functioning of the entire system Xi et al. (2023); Wang et al. (2023b). It has been extensively used in various fields such as biology An (2021), ecology González-Crespo et al. (2023), sociology Flache et al. (2022), etc., to model systems where individual entities influence collective behavior. However, the early attempts at agent-based simulation Farmer and Foley (2009); Geanakoplos (2010); Zheng et al. (2022) are limited because they are not autonomous and need human-defined rules or goals, which may not simulate the complex dynamics of urban systems. Hence, LLM-driven agents such as Urban Generative Intelligence (UGI) Xu et al. (2023b), can serve as up-to-date solutions for simulating urban dynamics based on cross-domain urban data. This paradigm not only propels forward the field of urban computing, but also paves the way for future cities that are more adaptive and responsive to the evolving needs of their inhabitants.
@@
  • Multi-modal Causal Learning: Causal learning or inference aims to investigate causal relationships between variables, ensuring stable and robust learning and inference Neuberg (2003); Pearl et al. (2016). Integrating deep learning techniques with causal inference has shown great success in recent years, especially in the fields of spatio-temporal graph forecasting Zhou et al. (2023b); Xia et al. (2024); Li et al. (2022a), CV Wang et al. (2020d); Zhang et al. (2020a); Lin et al. (2022), NLP Veitch et al. (2020); Zhang et al. (2023f); Tian et al. (2022), and recommender systems Zheng et al. (2021); Gao et al. (2022a). However, regarding cross-domain data fusion in urban computing, the application of causal learning is still in its early stages. Hence, we eagerly anticipate further research on multi-modal causal learning of urban data, aiming to improve the interpretability of intricate and dynamic urban systems.
@@
  • Multi-source Data Privacy: Urban data can be highly sensitive, especially in multi-source scenarios in the field of economy and healthcare. When models are trained upon such shared data, there is a risk that they may memorize specific information from the training data, potentially compromising the privacy of individuals. As a result, there is a need for further research on how to incorporate privacy-preserving techniques, such as differential privacy Dwork (2008); Yang et al. (2023b) and federated learning Zhang et al. (2021a); Li et al. (2021b), in order to safeguard data privacy while still benefiting from the advantages of cross-domain data fusion in urban computing.
@@
  • Computation Efficiency: Current urban research emphasizes the fulfillment of specific applications within cross-domain urban computing, but it overlooks the crucial aspect of computational efficiency. This oversight hampers the practical deployment of these computational models in real-life scenarios. Addressing this challenge requires our focus on optimizing computation efficiency, involving model compression (e.g., knowledge distillation Gou et al. (2021), low-rank decomposition Yao et al. (2023); Li et al. (2023e), and quantization Dettmers et al. (2022, 2023)), efficient training (e.g., prompt tuning Liu et al. (2023d) and hardware-assisted attention acceleration Dao et al. (2022); Dao (2023)), and efficient architecture (e.g., mixture of experts Chen et al. (2022c); Shen et al. (2023a)), to enhance the feasibility and effectiveness of deploying cross-domain data fusion solutions in practical urban environments.

### v2

@@ 5.6 Environment
Rapid urbanization may result in a potential threat to cities’ environment. For example, it is reported that ninety percent of air pollution is caused by urban transportation. Environment protection is an essential topic for urban computing. We have witnessed much research on the environment from different aspects of urban computing, such as air quality prediction [138, 114, 129, 279, 166], noise controlling [1, 117, 118, 64, 185] and weather forecasting [399, 341, 36, 247, 87]. Various data sources cloud provide information from various aspects for us to sense the environment. Further, the fusion those information from multiple sources was proven to be effective in understanding and predicting the variation of the environment in a city.
Zheng et al. [399] conducted research on cities’ air quality based on multi-source big data. They proposed a multi-view hybrid model to predict future 48-hour air quality of a single station point by converging historical air quality data of one station, current meteorological data in the area, weather forecasting data as well as air quality data from other stations. The various meteorological factors and spatial relations between different stations are both considered by effectively fusing these data which significantly improved the capability and accuracy of fine-grain air quality prediction. Further, Ma et al. [192] utilized a graph neural network to model spatio-temporal correlations between different meteorological variables and different stations as graph neural networks have demonstrated excellent performances for modeling spatio-temporal information. This kind of research makes city-level fine-grained weather forecasting possible with multi-source spatio-temporal data.
@@ 5.7.1 Energy Efficient Urban Planning
Electric vehicles are regarded as a promising solution for green transportation and sustainable cities. With the gradual expansion of the market share of electrical vehicles, the demand for car charging facilities is increasing. One important issue in urban planning is the selection of charging station locations. Proper selection of these locations not only optimizes city management but also meets the daily commuting needs of users, reducing unnecessary energy consumption during the search for charging stations. The selection of charging station locations should take into account various factors, such as the economic situation of the region, the number of electric vehicles in circulation, residents’ travel patterns, road conditions, and traffic conditions. This necessitates the integration and analysis of multiple data sources to estimate the regional charging demand and identify suitable locations for charging stations. Tu et al. [267] proposed a spatio-temporal demand coverage location model based on the taxi GPS data in Shenzhen, China as well as data from the charging stations. This model provided a charging station sitting approach by considering the waiting time which could hugely reduce energy consumption for charging. He et al. [100] developed a location method with the consideration of cars’ driving range extracted from the GPS data. Battaïa et al. [16] proposed a framework for charging situation sitting considering the real-life problem of charging infrastructure and sustainability.
@@ 5.7.2 Energy Efficient Urban Transportation
Transportation energy efficiency aims to recommend and navigate energy-efficient travel routes for every transportation participant. Nowadays, within the worldwide energy shortage, fuel prices have been continuously rising [387]. However, traffic congestion in 439 cities in the United States resulted in a total loss of 3.3 billion gallons of wasted fuel [237] in only one year. Traditional route planning studies primarily emphasized faster and more reliable routes, often overlooking the differences in energy consumption among different travel routes and modes for the same travel demand. For instance, in certain situations, taking a route through a congested city center may result in faster arrival times compared to taking a detour on a highway. Despite the first route being shorter in distance, the driver may experience delays due to frequent stops and slow speeds caused by city congestion.
In application, recommending energy-efficient routes for users is a highly complex task. Unlike traditional shortest path problems, it requires considering various factors such as vehicle efficiency, driving habits, road conditions, POIs, and more. The key to addressing this task lies in effectively integrating information from multiple sources. Wang et al. [293] researched personalized fuel-efficient route recommendations based on history trajectories. The temporal information from the trajectories as well as the driving factors are input into a transformer model to compose fuel efficiency prediction on the target route. By combining with a genetic algorithm for the best recommendation, this research could effectively find the fuel-efficient routes on the real-world dataset. DeepFEC [66] proposed a framework to predict road-level energy consumption through a fusion of contextual vehicle data such as type, weight and engine configuration, and spatio-temporal traffic data on roads. In this framework, a deep convolution residual model is designed to capture the spatial dynamic information from the data and a Bi-LSTM model is responsible for the temporal information. Oh et al. [209] is the first large dataset that combines trajectories of personal cars with users’ information, making it possible to mine users’ driving behaviors. Based on this multi-source dataset, Lai et al. [145] designed a meta-learning framework for predicting vehicle energy consumption in Ann Arbor, Michigan, USA that can be fine-tuned based on a user’s driving preference.
@@ 6 Challenges and Future Directions
While urban multi-modal research has made significant advancements in recent years, several challenging issues persist, highlighting directions for potential future research. As shown in Figure 27, we summarize these challenges and suggest potentially feasible research directions as follows:
Figure 27: Challenges and future directions of cross-domain data fusion in Urban Computing.
@@
  • LLM-enhanced Application: Since the advent of GPT-4 [211] and Sora [22], the academic community has embarked on extensive research on the roles of LLM as predictors [169, 106], enhancers [336, 221], controllers [73, 242, 317], and evaluators [255, 169]. However, the field of cross-domain data fusion in urban computing is still in its initial stage in terms of exploring how to apply LLMs effectively. For example, there exists research [232, 24, 195] focusing on LLM as urban predictors but also as naive reasoners yet; whereas others [105, 141] attempt to leverage LLM as model backbone but such works are limited within remote sensing domain instead of general urban computing. A critical limitation currently facing the application of LLMs in Urban Computing is their inability to perceive spatial relationships and dependencies inherent in natural spaces [408], a trait typically associated with language models. For instance, mainstream LLMs struggle to comprehend intricate spatial orientations and positions of urban entities. Moreover, the majority of publicly accessible LLMs are predominantly text-based and exhibit limited performance in other modalities. Nevertheless, we are encouraged by the growing number of successful initiatives in the graph learning domain that aim to enhance LLMs’ comprehension of complex relations within graph data [259, 160]. Furthermore, the latest multimodal LLMs, such as GPT-4o, demonstrate considerable promise in offering effective multimodal solutions. Therefore, we eagerly anticipate that the research community will focus on various applications of LLMs in urban computing, exploring their impact similar to their roles in NLP and other domains.
@@
  • Agent-based Simulation: Agent-based simulation aims to model the behavior of individual components or agents to comprehend their interactions and how they collectively contribute to the functioning of the entire system [310, 280]. It has been extensively used in various fields such as biology [6], ecology [89], sociology [72], etc., to model systems where individual entities influence collective behavior. However, the early attempts at agent-based simulation [70, 81, 389] are limited because they are not autonomous and need human-defined rules or goals, which may not simulate the complex dynamics of urban systems. Hence, LLM-driven agents such as Urban Generative Intelligence (UGI) [317], can serve as up-to-date solutions for simulating urban dynamics based on cross-domain urban data. This paradigm not only propels forward the field of urban computing, but also paves the way for future cities that are more adaptive and responsive to the evolving needs of their inhabitants.
@@
  • Multi-modal Causal Learning: Causal learning or inference aims to investigate causal relationships between variables, ensuring stable and robust learning and inference [206, 220]. Integrating deep learning techniques with causal inference has shown great success in recent years, especially in the fields of spatio-temporal graph forecasting [403, 312, 149], CV [290, 359, 168], NLP [269, 371, 264], and recommender systems [393, 75]. However, regarding cross-domain data fusion in urban computing, the application of causal learning is still in its early stages. One of the most severe challenges is to represent the complex cross-modal causality in urban scenarios. The deep learning community has yet to converge on a universally accepted approach for accurately yet efficiently representing cross-modal relations. However, as an increasing number of studies delve into the semantic essence of information across various modalities and advance the fields of graph learning, representation learning, and urban foundational models, we eagerly anticipate further research on multi-modal causal learning of urban data, aiming to improve the interpretability of intricate and dynamic urban systems.
@@
  • Multi-source Data Privacy: Urban data can be highly sensitive, especially in multi-source scenarios in the field of economy and healthcare. When models are trained upon such shared data, there is a risk that they may memorize specific information from the training data, potentially compromising the privacy of individuals. Particularly in the context of deploying data fusion models, the inclusion of data from diverse sources and modalities can indeed bolster the models’ performance. However, this also heightens the risk of inadvertently exposing sensitive privacy data. Given these privacy concerns, many institutions and city governments are reluctant to share urban data, thereby impeding the deployment of deep learning models. This poses a significant barrier to the advancement of the community. Consequently, there is a pressing need for research that explores the integration of privacy-preserving techniques, such as differential privacy [65, 326] and federated learning [357, 153]. The goal is to protect data privacy while still reaping the benefits of cross-domain data fusion in urban computing.
@@
  • Computation Efficiency: Current urban research emphasizes the fulfillment of specific applications within cross-domain urban computing, but it overlooks the crucial aspect of computational efficiency. This oversight hampers the practical deployment of these computational models in real-life scenarios. Addressing this challenge requires our focus on optimizing computation efficiency, involving model compression (e.g., knowledge distillation [90], low-rank decomposition [331, 159], and quantization [58, 59]), efficient training (e.g., prompt tuning [180] and hardware-assisted attention acceleration [55, 54]), and efficient architecture (e.g., mixture of experts [44, 241]), to enhance the feasibility and effectiveness of deploying cross-domain data fusion solutions in practical urban environments.